# A Prefrontal Cortex-inspired Architecture for Planning in Large Language Models

### v3

@@ 2.4 Plan Generation
Algorithm 3 describes the complete LLM-PFC algorithm. To generate a plan, a set of subgoals is first generated by the $\operatorname{TaskDecomposer}$ based on the final goal and current state. These subgoals are then pursued one at a time, utilizing the $\operatorname{Search}$ loop to generate actions until the $\operatorname{Orchestrator}$ determines that the subgoal has been achieved. The actions are accumulated in a plan buffer $P$ until either the $\operatorname{Orchestrator}$ determines that the final goal has been reached, or the maximum allowable number of actions $T$ are accumulated. This approach is inspired by the role that aPFC plays in task decomposition. This involves the decomposition of tasks into smaller, more manageable tasks, and the coordinated sequential execution of these component tasks (Ramnani & Owen, 2004).
@@ 3 Experiments
Experiment details are described in Section A.2. Code is available at: https://github.com/Shanka123/llm_pfc.
@@ 3.1 Tasks
Graph Traversal. We performed controlled experiments on four multi-step planning tasks based on graph traversal using the CogEval protocol (Momennejad et al., 2023). Natural language descriptions of a graph are provided with each node assigned to a room (e.g., ‘room 4 is connected to room 7’). We focused on a particular type of graph (Figure 4) with community structure (Schapiro et al., 2013) previously found to be challenging for a wide variety of LLMs. The first task, Valuepath, involves finding the shortest path from a given room that results in the largest reward possible. A smaller reward and a larger reward are located at two different positions in the graph. We fixed the two reward locations, and created 13 problems based on different starting locations. The second task, Steppath, involves finding the shortest path between a pair of nodes. We evaluated problems with an optimal shortest path of 2, 3, or 4 steps. We generated 20 problems for each of these conditions by sampling different starting and target locations.
Tower of Hanoi. We also investigated a classic multi-step planning task called the Tower of Hanoi (ToH) (Figure 5). In the original formulation, there are three pegs and a set of disks of different sizes. The disks are stacked in order of decreasing size on the leftmost peg. The goal is to move all disks to the rightmost peg, such that the disks are stacked in order of decreasing size. There are a couple of rules that determine which moves are considered valid. First, a disk can only be moved if it is at the top of its stack. Second, a disk can only be moved to the top of another stack if it is smaller than the disks in that stack (or if the peg is empty). More complex versions of the task can be created by using a larger number of disks.
Logistics. To assess the ability to generate plans in more real-world settings, we investigated a logistics plan generation task involving the transportation of goods between cities using airplanes and trucks (more details can be found in Valmeekam et al. (2023)).
Figure 2: Graph traversal results. Top row: Valuepath results. Middle row: Steppath results. Bottom row: Detour and Reward Revaluation results. ‘solved’ indicates percentage of problems solved without proposing invalid actions ($\uparrow$ better). ‘invalid’ indicates percentage of moves that are invalid ($\downarrow$ better). ‘Plan steps’ indicates number of steps in plan for solved problems only (therefore excluding many problems for the baseline models; $\downarrow$ better). GPT-4 Zero-shot and ICL baselines are deterministic, and therefore a single run was performed on all problems. Note that LLM-PFC did not employ tree search on the Steppath task, and did not employ task decomposition on any of the graph traversal tasks. Without tree search, LLM-PFC’s performance is deterministic, and therefore only a single run was performed on the Steppath task. Gray error bars reflect 95% binomial confidence intervals (for models evaluated on a single run). Dots reflect values of 0%. For Valuepath, we performed 10 runs with LLM-PFC, and present average performance $\pm$ the standard error of the mean (black error bars). ICL: In Context Learning. PFC: Prefrontal Cortex.
@@ 3.2 Baselines
We compared our model to several baseline methods. The first method involved asking GPT-4 (zero-shot) to provide the solution step by step. For the second method, in-context learning (ICL), we provided GPT-4 with a few in-context examples of a complete solution. We provided two examples for ToH and Valuepath, and 3 examples (one each for 2, 3, and 4 steps) for Steppath. The third method was chain-of-thought (CoT) (Wei et al., 2022b). For this method, the in-context examples were annotated with a series of intermediate computations that break down the planning process into multiple steps (see Sections A.6.7-A.6.9 for example baseline prompts). Finally, we evaluated multi-agent debate (MAD), using the codebase from Du et al. (2023). In this approach, similar to LLM-PFC, a solution is generated through the interaction between multiple LLM instances (each instance was equivalent to the GPT-4 ICL baseline); however, unlike LLM-PFC, these instances are not specialized to perform specific functions.
Figure 3: Tower of Hanoi (ToH) results. ‘solved’ indicates percentage of problems solved without proposing invalid actions ($\uparrow$ better). ‘invalid’ indicates percentage of moves that are invalid ($\downarrow$ better). Note that 4-disk problems are out-of-distribution (OOD). GPT-4 Zero-shot, ICL, and CoT ICL baselines are deterministic and reflect a single run. Gray error bars reflect 95% binomial confidence intervals. Dots reflect values of 0%. LLM-PFC results for 3-disk problems reflect the average over 5 runs $\pm$ the standard error of the mean (black error bars). LLM-PFC results for 4-disk problems reflect a single run, due to the high computational cost of multiple runs. ICL: In Context Learning. CoT: Chain of Thought. MAD: Multi-Agent Debate. PFC: Prefrontal Cortex.
@@ 4 Results
Figure 2 shows the results on the four graph traversal tasks (see Section A.4 for all results in Table form). On the Valuepath task, LLM-PFC solved 100% of problems, significantly outperforming both baselines. On the Steppath task, LLM-PFC displayed perfect performance for 2-step and 3-step paths, and near-perfect performance for 4-step paths, again significantly outperforming both baselines. Notably, LLM-PFC’s proposed plans were close to the optimal number of steps for both tasks. LLM-PFC also significantly outperformed both baselines on the Detour and Reward Revaluation tasks, with near-perfect performance in the Detour task. This demonstrates that LLM-PFC can flexibly adjust to new circumstances when generating plans. Finally, the model did not propose any invalid actions in any of the four tasks (e.g., it did not hallucinate the presence of non-existent edges), due to the filtering of invalid actions by the $\operatorname{Monitor}$.
@@
Figure 3 shows the results on Tower of Hanoi (ToH). LLM-PFC demonstrated a significant improvement both in terms of the number of problems solved (left) and the number of invalid actions proposed (right). On 3-disk problems, LLM-PFC yielded a nearly seven-fold improvement in the number of problems solved over zero-shot performance, and also significantly outperformed standard in-context learning (ICL), chain-of-thought (CoT ICL), and multi-agent debate (MAD). LLM-PFC’s improved performance relative to MAD (Du et al., 2023) demonstrates the importance of interaction between specialized LLM instances (i.e., a modular approach), whereas MAD involves interactions between multiple LLM instances prompted to perform the same task. For the problems that LLM-PFC solved, the average plan length (5.4) was close to the optimal number of moves (4.4). The model also demonstrated some ability to generalize out-of-distribution (OOD) to more complex 4-disk problems (not observed in any in-context examples), whereas the baseline models solved close to 0% of these problems. Notably, LLM-PFC did not propose any invalid actions, even on OOD 4-disk problems, whereas the baselines proposed a significant number of invalid actions.
Finally, we found that LLM-PFC also significantly improved performance on the logistics task, successfully solving 31% (62/200) of problems [^1] Note that we did not use tree search or the $\operatorname{TaskDecomposer}$ on these problems. Incorporating these components may further improve the performance of LLM-PFC on this task., whereas the GPT-4 ICL baseline solved only 10.5% (21/200) of problems, and GPT-4 zero-shot solved only 7.5% (15/200) of problems (as reported in the original paper (Valmeekam et al., 2023)). This demonstrates the potential of LLM-PFC to be beneficial in more real-world planning domains.
@@ 4.1 Ablation Study
We also carried out an ablation study to determine the relative importance of each of LLM-PFC’s major components, focusing on the 3-disk ToH problems. Figure 3 (left) shows the results. We found that the $\operatorname{Monitor}$ was the most important component, as ablating this module resulted in significantly fewer solved problems, due primarily to an increased tendency to propose invalid moves (31% invalid moves vs. 0% for other ablation models). Ablating the tree search and $\operatorname{TaskDecomposer}$ module also resulted in significantly fewer solved problems. Overall, these results suggest that all major components played an important role in the model’s performance. These results also demonstrate that LLM-PFC’s improved performance was not due entirely to the use of tree search (which is shared with tree-of-thoughts (Yao et al., 2023)), but also depended on the incorporation of other modules, including the $\operatorname{TaskDecomposer}$, and especially the $\operatorname{Monitor}$.
@@ 5 Related Work
A few recent studies have investigated planning in LLMs. These studies suggest that, although LLMs can perform relatively simple planning tasks (Huang et al., 2022), and can learn to make more complex plans given extensive domain-specific fine-tuning (Pallagani et al., 2022; Wu et al., 2023), they struggle on tasks that require zero-shot or few-shot generation of complex multi-step plans (Valmeekam et al., 2023; Momennejad et al., 2023). These results also align with studies that have found poor performance in tasks that involve other forms of extended multi-step reasoning, such as arithmetic (Dziri et al., 2023). Our approach is in large part motivated by the poor planning and reasoning performance exhibited by LLMs in these settings.
@@
Some recent approaches have employed various forms of heuristic search to improve performance in LLMs (Lu et al., 2021; Zhang et al., 2023), but these approaches have generally involved search at the level of individual tokens. This is in contrast to our approach, in which search is performed at the more abstract level of task states (described in natural language). This is similar to other recently proposed black-box approaches in which ‘thoughts’ – meaningful chunks of natural language – are utilized as intermediate computations to solve more complex problems. These approaches include scratchpads (Nye et al., 2021), chain-of-thought (Wei et al., 2022b), tree-of-thoughts (Yao et al., 2023), reflexion (Shinn et al., 2023), Society of Mind (Du et al., 2023), and Describe-Explain-Plan-Select (Wang et al., 2023). All of these approaches can be viewed as implementing a form of controlled, or ‘system 2’, processing (as contrasted with automatic, or ‘system 1’, processing) (Schneider & Shiffrin, 1977; Sloman, 1996; Kahneman, 2011). In the brain, these controlled processes are strongly associated with the prefrontal cortex (Miller & Cohen, 2001). Therefore, in the present work, we leveraged knowledge from cognitive neuroscience about the modular properties of the PFC. The resulting architecture shares some components with other black box approaches (e.g., tree search (Yao et al., 2023)), but also introduces a number of new components (error monitoring, task decomposition, task coordination, state/action distinction), and combines these components in a novel manner inspired by the functional organization of the human brain (see Section A.1).
@@
There have also been a number of proposals for incorporating modularity into deep learning systems, including neural module networks (Andreas et al., 2016), and recurrent independent mechanisms (Goyal et al., 2019). Our approach is distinguished from these approaches by the proposal of modules that perform specific high-level component processes, based on knowledge of specific subregions within the PFC. Finally, our approach is closely related to a recent proposal to augment deep learning systems with PFC-inspired mechanisms (Russin et al., 2020). LLM-PFC can be viewed as a concrete framework for accomplishing this goal.

### v4

@@ 2.4 Plan Generation
Algorithm 1 describes the complete MAP algorithm. To generate a plan, the $\operatorname{TaskDecomposer}$ component of MAP first generates a set of subgoals based on the final goal and current state. These subgoals guide the search and are internally pursued one at a time, utilizing the $\operatorname{Search}$ loop to generate actions until the $\operatorname{Orchestrator}$ determines that the subgoal has been achieved. The actions are accumulated in a plan buffer $P$ until either the $\operatorname{Orchestrator}$ determines that the final goal has been reached, or the maximum allowable number of actions $T$ are accumulated.
@@ 3 Experiments
Experiment details are described in Section A.3. Code is available at: https://github.com/MAPLLM/MAPICLR2025sub.
@@ 3.1 Tasks
Graph Traversal. We performed experiments on four multi-step planning tasks based on graph traversal using the CogEval protocol (Momennejad et al., 2023). Natural language descriptions of a graph are provided with each node assigned to a room (e.g., ‘room 4 is connected to room 7’). The tasks included Valuepath, involving finding the shortest path between a given room and the largest of two possible rewards (but without going through the room with the smaller reward); Steppath, involving finding the shortest path between two rooms; Detour, in which the Valuepath task is first described, after which an edge is subsequently removed from the graph; and Reward Revaluation, in which the Valuepath task is first described, and the value associated with two reward locations is subsequently changed. Please see Section A.4 in the Appendix for more details.
Tower of Hanoi. We also investigated a classic multi-step planning task called the Tower of Hanoi (ToH) (Figure 5). In the original task, there are three pegs and a set of disks of different sizes. The disks must be moved into a particular goal configuration, while observing a set of constraints that prevent simple solutions. In our experiments, we designed an alternative (but isomorphic) formulation of this task in which the inputs are text-based rather than visual. This text-based formulation made it possible to evaluate language models on the task, but it also resulted in a task that does not share any surface features with the original task, making it unlikely that GPT-4 could rely on exposure to descriptions of ToH in its training data to solve the problem. Please see Section A.4 in the Appendix for more details.
PlanBench. To assess the generality and robustness of our approach, we also investigated a more extensive planning benchmark, PlanBench, consisting of synthetically generated problems in a number of distinct domains. We specifically investigated the Logistics domain, involving the transportation of goods between cities using airplanes and trucks, and the Mystery Blocksworld (deceptive) domain, which involves arbitrary names for entities and actions, and is the most challenging domain in the dataset (more details can be found in Valmeekam et al. (2023)).
StrategyQA. Finally, to test the extent to which MAP can be applied to more real-world tasks, we investigated StrategyQA, an NLP task that requires multi-step reasoning, and has proven challenging for standard LLM methods (Geva et al., 2021). In this task, an unusual question is posed (e.g. ‘Did Aristotle own a laptop?’) that requires multi-step reasoning. The question must be decomposed into sub-questions which must be successively solved in order to arrive at a final answer (more details can be found in Geva et al. (2021))
@@ 3.2 Baselines
We compared our model to several baseline methods. The first method involved asking GPT-4 (zero-shot) to provide the solution step by step. For the second method, in-context learning (ICL), we provided GPT-4 with a few in-context examples of a complete solution. We provided two examples for ToH, Valuepath, Detour, and Reward Reval, and three examples for Steppath (one each for 2, 3, and 4 steps) and PlanBench. The third method was chain-of-thought (CoT) (Wei et al., 2022b). For this method, the in-context examples were annotated with a series of intermediate computations that break down the planning process into multiple steps (see Sections A.7.7-A.7.9 for example baseline prompts). The fourth method was multi-agent debate (MAD), using the codebase from Du et al. (2023). In this approach, similar to MAP, a solution is generated through the interaction between multiple LLM instances (each instance was equivalent to the GPT-4 ICL baseline); however, unlike MAP, these instances are not specialized to perform specific functions. Finally, we investigated tree-of-thought (ToT), using the original codebase from  Yao et al. (2023). Similar to MAP, ToT uses multiple LLM modules to perform tree search, although MAP incorporates additional modules and control processes (see Section A.2). To ensure that ToT was given the best chance of performing well on our tasks, we tested two versions, one with prompts that were similar to those in the original implementation (shown in the main results Section 4), and one that incorporated prompts from MAP (ToT-MAP, see Appendix Table 8). For ToT, multiple potential plans are generated for each problem, and a method is required to select one of these plans (a problem-specific heuristic was used in the original work). To ensure a fair comparison, we evaluated using two metrics, 1) the best plan (according to a groundtruth evaluation) for each problem, 2) the average performance of all plans for each problem. We report both of these metrics for both ToT and MAP.
@@ 4 Results
Figure 2: Graph traversal results. ‘solved’ indicates percentage of problems solved without proposing invalid actions ($\uparrow$ better). GPT-4 Zero-shot, ICL, COT, and MAD baselines are deterministic, and therefore a single run was performed on all problems. Note that MAP did not employ tree search on the Steppath task, and did not employ task decomposition on any of the graph traversal tasks. Without tree search, MAP’s performance is deterministic, and therefore only a single run was performed on the Steppath task, whereas we performed 5 runs with ToT. Gray error bars reflect 95% binomial confidence intervals (for models evaluated on a single run). Dots reflect values of 0%. Dark bars indicate average performance over multiple plans/runs. Light bars indicate best performance. For Valuepath, Detour, and Reward Revaluation we performed 10, 10, and 5 runs respectively with MAP and ToT, and present average performance $\pm$ the standard error of the mean (black error bars).
Figure 3: Tower of Hanoi (ToH) results. ‘solved’ indicates percentage of problems solved without proposing invalid actions ($\uparrow$ better). ‘invalid’ indicates percentage of moves that are invalid ($\downarrow$ better). Note that 4-disk problems are out-of-distribution (OOD). GPT-4 Zero-shot, ICL, CoT, and MAD baselines are deterministic and reflect a single run. Gray error bars reflect 95% binomial confidence intervals. Dots reflect values of 0%. Dark bars indicate average performance over multiple plans/runs. Light bars indicate best performance. MAP results for 3-disk problems reflect the average over 5 runs $\pm$ the standard error of the mean (black error bars). MAP results for 4-disk problems reflect a single run, due to the high computational cost of multiple runs.
@@
Figure 2 shows the results on the four graph traversal tasks (see Section A.5 for all results in tabular form). On the Valuepath task, MAP solved 100% of problems, outperforming all baselines. On the Steppath task, MAP displayed perfect performance for 2-step and 3-step paths, and near-perfect performance for 4-step paths, again outperforming all baselines. MAP also outperformed all baselines on the Detour task, and performed on par with the baselines on the Reward Revaluation task (while still outperforming GPT-4 zero-shot). This demonstrates that MAP can flexibly adjust to new circumstances when generating plans. Finally, the model did not propose any invalid actions in any of the four tasks (i.e., it did not hallucinate the presence of non-existent edges), due to the filtering of invalid actions by the $\operatorname{Monitor}$ (Figure 6).
@@
Figure 3 shows the results on Tower of Hanoi (ToH). MAP demonstrated a significant improvement both in terms of the number of problems solved (left) and the number of invalid actions proposed (right). On 3-disk problems, MAP yielded a nearly seven-fold improvement in the number of problems solved over zero-shot performance, and significantly outperformed standard in-context learning (ICL), chain-of-thought (CoT), multi-agent debate (MAD), and tree-of-thought (ToT). When considering the best plan (out of 5 runs) for each problem, MAP achieved a perfect score of 100%. MAP’s improved performance relative to MAD demonstrates the importance of interaction between specialized LLM instances (i.e., a modular approach), whereas MAD involves interactions between multiple LLM instances prompted to perform the same task. MAP’s superior performance relative to ToT demonstrates that tree search, though an important part of the approach, is not sufficient to explain MAP’s performance, and the other modules play an important role. For the problems that MAP solved, the average plan length (5.4) was close to the optimal number of moves (4.4). The model also demonstrated some ability to generalize out-of-distribution (OOD) to more complex 4-disk problems (not observed in any in-context examples), whereas the baseline models solved close to 0% of these problems. Notably, MAP did not propose any invalid actions, even on OOD 4-disk problems, whereas the baselines proposed a significant number of invalid actions.
We also performed several experiments with ToH to better understand the MAP algorithm. First, we investigated how MAP’s performance varied as a function of the depth of tree search. We found that a depth of $L=2$ provided the best tradeoff between performance and cost (Table 14). A depth of $L=1$ resulted in worse performance, while a depth of $L=3$ incurred greater cost without significantly improving performance. Second, we investigated the extent to which MAP’s performance on this task depended on being provided with an explicit strategy for task decomposition (the goal recursion strategy). To address this, we provided the GPT-4 zero-shot, ICL, and CoT baselines with a description of this strategy. We found that this strategy did improve baseline performance, but MAP still outperformed these baselines even when they were provided with the strategy (Table 12). Third, we investigated whether the computational costs of the MAP algorithm could be mitigated by using a smaller LLM. We found that a version of MAP that used Llama3-70B still outperformed baselines that used the same LLM, and even outperformed the best GPT-4 baseline, GPT-4 ICL (Table 10).
Table $\dagger$2 shows the results for the PlanBench dataset, where MAP outperformed all of the baselines that we considered. Notably, due to the complexity of the problems in this dataset, it was very costly to perform tree search, so we evaluated a minimal version of MAP that did not involve tree search. For this same reason, we were unable to evaluate a ToT baseline on the full set of problems, but we include a comparison with ToT on a subset of problems in Table 13. Even without tree search, MAP outperformed ToT. These results demonstrate that, although MAP certainly benefits from the use of tree search, it can still provide significant performance benefits in domains where this is not feasible. We also investigated a zero-shot version of MAP (without in-context examples) on the most challenging PlanBench domain (mystery blocksworld), and found that it outperformed both GPT-4 zero-shot and GPT-4 ICL (Table 11), suggesting that MAP can be useful even in settings where in-context examples are not available.
Table $\dagger$2 shows the results for the StrategyQA benchmark, where MAP outperformed both CoT and ToT, and performed on par with human participants. This demonstrates the potential of MAP to be beneficial in more real-world tasks, such as question-answering tasks that require multi-step reasoning.
[…]
Model Logistics Mystery BW
GPT-4 Zero-shot 7 0.2
GPT-4 ICL 12 7.8
MAD 16.2 7.3
GPT-4 CoT 17 10.6
MAP 24 27.4
[…]
Model Accuracy
[…]
GPT-4 CoT 84.7 $\pm$ 0.3
MAP 87.7 $\pm$ 0.7
[…]
Finally, we performed transfer experiments to study whether few-shot in-context learning would support generalization to different planning tasks. Table 3 shows the results for these experiments, including results for transfer from planning on a smaller graph to planning on a larger graph (n7tree $\rightarrow$ n15star), transfer to a semantically distinct but structurally isomorphic task (blocksworld (BW) $\rightarrow$ mystery blocksworld (mystery BW)), and transfer between completely different tasks (ToH $\rightarrow$ Mystery BW). We found that MAP outperformed both GPT-4 ICL and CoT in each of these settings, indicating that MAP can improve the generalizability and robustness of planning in LLMs.
Table 3: Transfer between different planning tasks. Results reflect % solved problems.
Model n7tree $\rightarrow$ n15star Valuepath BW $\rightarrow$ Mystery BW ToH $\rightarrow$ Mystery BW
GPT-4 ICL 51 0.2 0
GPT-4 CoT 65 1.4 0
MAP 80 12.2 6.6
@@ 4.1 Ablation study
We also carried out an ablation study to determine the relative importance of each of MAP’s major components, focusing on the 3-disk ToH problems. Figure 3 (left) shows the results. We found that the $\operatorname{Monitor}$ was the most important component, as ablating this module resulted in significantly fewer solved problems, due primarily to an increased tendency to propose invalid moves (31% invalid moves vs. 0% for other ablation models). This highlights the importance of having a separate, modularized monitoring process. Ablating the tree search and $\operatorname{TaskDecomposer}$ module also resulted in significantly fewer solved problems. The impaired performance following the ablation of the tree search indicates the benefit of considering the multi-step implications of proposed actions, rather than committing to a single action as is done in standard autoregressive methods such as chain-of-thought. The impaired performance following the ablation of the $\operatorname{TaskDecomposer}$ highlights the benefit of decomposing a task into subgoals, which allows MAP to factorize a complex task into a set of smaller, more manageable tasks. Overall, these results suggest that all major components played an important role in MAP’s performance. Moreover, the improved performance was not due entirely to the use of tree search (which is shared with tree-of-thoughts (Yao et al., 2023)), but also depended on the incorporation of other modules such as the $\operatorname{TaskDecomposer}$ and especially the $\operatorname{Monitor}$.
@@ 5 Related work
A few recent studies have investigated planning and multi-step decision making in LLMs. These studies suggest that, although LLMs can perform relatively simple planning tasks (Huang et al., 2022), and can learn to make more complex plans given extensive domain-specific fine-tuning (Pallagani et al., 2022; Wu et al., 2023), they struggle on tasks that require zero-shot or few-shot generation of multi-step plans (Valmeekam et al., 2023; Momennejad et al., 2023). These results also align with studies that have found poor performance in tasks that involve other forms of extended multi-step reasoning, such as arithmetic (Dziri et al., 2023). Our approach is in large part motivated by the poor planning and reasoning performance exhibited by LLMs in these settings.
@@
Some recent approaches have employed various forms of heuristic search to improve performance in LLMs (Lu et al., 2021; Zhang et al., 2023b), but these approaches have generally involved search at the level of individual tokens. Importantly, this is in contrast to our approach, in which search is performed at the more abstract level of task states (described in natural language). Ours is similar to other recently proposed black-box approaches in which ‘thoughts’ – meaningful chunks of natural language – are utilized as intermediate computations to solve more complex problems. These approaches include scratchpads (Nye et al., 2021), chain-of-thought (Wei et al., 2022b), tree-of-thoughts (Yao et al., 2023), reflexion (Shinn et al., 2023), multi-agent methods (Du et al., 2023; Zhang et al., 2023a), Describe-Explain-Plan-Select (Wang et al., 2023), and methods for combining planning with external tools (Ruan et al., 2023; Kong et al., 2023). All of these approaches can be viewed as implementing a form of controlled, or ‘system 2’, processing (as contrasted with automatic, or ‘system 1’, processing) (Schneider & Shiffrin, 1977; Sloman, 1996; Kahneman, 2011). Our approach has a similar high-level motivation, and shares some components with other black box approaches (e.g., tree search (Yao et al., 2023)), but also introduces a number of new components (error monitoring, task decomposition, task coordination, state/action distinction), and combines these components in a novel manner (see Section A.2 for further discussion).
@@
There have also been a number of proposals for incorporating modularity into deep learning systems, including neural module networks (Andreas et al., 2016), and recurrent independent mechanisms (Goyal et al., 2019). Ours is distinguished from these approaches by the use of black-box modules that perform specific high-level functions (many of which are inspired by formal theories of decision-making, as discussed below), rather than merely incorporating a general bias toward modularity.
Our approach is inspired by formal theories of decision-making and planning, and has a particularly close connection to reinforcement learning (RL) (Sutton & Barto, 2018). In particular, many of the modules in the MAP algorithm are closely related to aspects of traditional RL algorithms. Specifically, the $\operatorname{Actor}$ and $\operatorname{Evaluator}$ modules bear some resemblance to the actor and the critic in the popular actor-critic framework (Barto et al., 1983), and the TaskDecomposer is related to hierarchical RL (Sutton et al., 1999; Dietterich, 2000; Bacon et al., 2017), in which temporal abstractions are learned to achieve subgoals. The $\operatorname{Predictor}$ is also closely related to the world model that can substitute for direct interation with the environment in model-based RL (Sutton & Barto, 2018; Daw, 2012). An important difference in each of these cases is that the modules in MAP only receive a task description and a couple of examples, relying on the general-purpose knowledge of the LLM to effectively perform the task, rather than being trained through RL. This also distinguishes the approach from other recent efforts to combine LLMs and RL (Carta et al., 2023; Zhou et al., 2024; Zhai et al., 2024), which involve training with RL through direct interaction with an external environment, whereas MAP generates plans internally. Finally, there are also some modules that have no obvious analog in previous RL algorithms, but which were necessitated by weaknesses that LLMs display in the planning domain. These include the $\operatorname{Monitor}$, which was necessitated by the tendency of LLMs to hallucinate or violate task constraints, and the $\operatorname{Orchestrator}$, which allows MAP to autonomously determine when a goal has been achieved (without groundtruth evaluation) and thus terminate planning.